### v2

# TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design

Abstract
Generalizing deep reinforcement learning agents to unseen environments remains a significant challenge. One promising solution is Unsupervised Environment Design (UED), a co-evolutionary framework in which a teacher adaptively generates tasks with high learning potential, while a student learns a robust policy from this evolving curriculum. Existing UED methods typically measure learning potential via regret, the gap between optimal and current performance, approximated solely by value-function loss. Building on these approaches, we introduce the transition prediction error as an additional term in our regret approximation. To capture how training on one task affects performance on others, we further propose a lightweight metric called co-learnability. By combining these two measures, we present Transition-aware Regret Approximation with Co-learnability for Environment Design (TRACED). Empirical evaluations show that TRACED yields curricula that improve zero-shot generalization across multiple benchmarks while requiring up to 2$\times$ fewer environment interactions than strong baselines. Ablation studies confirm that the transition prediction error drives rapid complexity ramp-up and that co-learnability delivers additional gains when paired with the transition prediction error. These results demonstrate how refined regret approximation and explicit modeling of task relationships can be leveraged for sample-efficient curriculum design in UED.
1 Introduction
Deep reinforcement learning (RL) has achieved remarkable success in games, continuous control, and robotics [1]. Ideally, we want agents that generalize robustly to a broad range of unseen environments. However, hand-crafting a training distribution that captures all real-world variability is intractable, and agents often overfit even large training sets, performing poorly out-of-distribution [2, 3].
Unsupervised Environment Design (UED) tackles this by adapting the curriculum: a teacher module generates training tasks that challenge the student agent [4]. A popular class of UED methods measures task difficulty by regret, the difference between the optimal return and the agent’s achieved return, and uses this metric to guide curriculum design [4, 5, 6, 7, 8, 9, 10]. Unfortunately, genuine regret requires knowing each environment’s optimal $Q^{*}$, which is infeasible in complex domains. Existing approaches, therefore, resort to coarse proxies such as Positive Value Loss (PVL) or maximum observed return (MaxMC) [11]. In this paper, we refine regret estimation by explicitly decomposing it into (i) a value error, captured by PVL, and (ii) a transition prediction error, estimated via a transition dynamics model.
We further introduce a metric called Co-Learnability to quantify how training on one task benefits others. For instance, consider three 100-word vocabulary tasks, Spanish, English, and Japanese, whose transfer patterns differ: because Spanish and English share many cognates, learning Spanish accelerates lexical access and boosts English accuracy [12, 13, 14], reflecting high Co-Learnability; in contrast, Japanese is typologically distant [15], so gains from Japanese may transfer unefficiently to English, reflecting low Co-Learnability. We present a lightweight estimator of Co-Learnability that leverages observed changes in approximated regret, avoiding any additional modeling overhead in the UED loop.
Figure 1: Task Priority Landscape. Task in the top-right quadrant (high difficulty and high co-learnability) are scheduled with the highest priority in the curriculum.
As illustrated in Figure 1, these two ingredients, refined regret and Co-Learnability, are fused into a unified Task Priority score via a rank-based transform (cf. [7]), which governs both new task generation and prioritized replay. We call our algorithm TRACED (Transition-aware Regret Approximation with Co-learnability for Environment Design). We evaluate TRACED on procedurally generated MiniGrid (MG) [16] and BipedalWalker (BW) [17]. In MG, we compare against Domain Randomization (DR) [18], PLR [5], ADD [19], and ACCEL [7], the strongest of these baselines. In BW, we additionally include the state-of-the-art (sota) CENIE [20].
On 12 held-out MG mazes, TRACED achieves a mean solved rate of 66% after 10k updates, outperforming ACCEL’s 60% after 20k updates. In the extreme maze variants PerfectMaze, TRACED reaches 27% success on PerfectMazeLarge and 10% on PerfectMazeXL, representing 35-40% relative gains over ACCEL with only half the training time. In BW, TRACED achieves an 18% improvement over CENIE while halving the wall-clock time. Ablation studies reveal that the transition-prediction error term drives rapid complexity ramp-up, and that Co-Learnability provides additional gains when combined with our regret term. Collectively, these results pave a new path toward sample-efficient curriculum design in UED by combining refined regret approximation with explicit modeling of task relationships.
We provide the full implementation of TRACED to facilitate the reproduction of our main results at https://github.com/Cho-Geonwoo/TRACED.
2 Preliminaries
2.1 Underspecified Partially Observable MDPs (UPOMDPs)
We model our environments as underspecified partially observable Markov decision processes (UPOMDPs) following Dennis et al. [4]. A UPOMDP is a tuple $\mathcal{M}=\langle A,O,\Theta,S,P_{0},P_{T},\mathcal{I},\mathcal{R},\gamma\rangle$, where $A$ is a finite set of actions, $S$ is a latent state space, and $O$ is an observation space. The observation function $\mathcal{I}:S\to\Delta(O)$ generates each observation $o_{t}$ given the true state $s_{t}$, the reward function $\mathcal{R}:S\times A\to\mathbb{R}$ and discount factor $\gamma\in[0,1)$ are shared across all levels. Crucially, $\Theta$ is a set of underspecified parameters that distinguish individual “levels”: for each $\theta\in\Theta$, the initial state is drawn from $P_{0}(\theta)\in\Delta(S)$ and transitions follow $P_{T}(s_{t+1}\mid s_{t},a_{t},\theta)$.
At each time step $t$, the agent observes $o_{t}\sim\mathcal{I}(s_{t})$ and selects an action $a_{t}$ according to a trajectory-conditioned policy $\pi\bigl{(}a_{t}|o_{0},a_{0},\dots,o_{t}\bigr{)}$. For a fixed level $\theta$, the utility of policy $\pi$ is the expected discounted return $U_{\theta}(\pi)=\mathbb{E}\!\Bigl{[}\sum_{t=0}^{T}\gamma^{t}\,r_{t}\Bigr{]},r_{t}=\mathcal{R}(s_{t},a_{t})$ with the expectation taken over both the stochastic dynamics and the policy’s choices. We denote an optimal policy on level $\theta$ by $\pi^{\star}_{\theta}\in\arg\max_{\pi}U_{\theta}(\pi)$.
2.2 Unsupervised Environment Design (UED)
UED provides a series of levels with unknown parameters that are used to produce task environments automatically as a curriculum for the agent so as to efficiently train a single generalist policy $\pi_{\phi}$ across the entire parameter space $\Theta$ [4]. Recent UED methods maximize regret of the agent to generate a distribution of environments that guide effective learning [5, 7, 20]. The regret of policy $\pi$ is defined as the difference in expected reward between optimal policy $\pi^{*}$ and policy $\pi$, $Regret_{\theta}(\pi)=U_{\theta}(\pi^{*})-U_{\theta}(\pi)$.
Because the true optimal policy $\pi^{\star}$ is unknown in complex environments, existing methods approximate the instantaneous regret using proxy metrics. For example, PLR [6] evaluates two such proxies: Positive Value Loss (PVL) and Maximum Monte Carlo (MaxMC). PVL estimates the instantaneous regret as the average positive part of the Generalized Advantage Estimation (GAE)-based Temporal Difference (TD) errors over an episode. PVL for an episode $\tau$ of length $T$ is defined as
$\mathrm{PVL}(\tau)=\frac{1}{T}\sum_{t=0}^{T}\max({\sum_{k=t}^{T}(\gamma\lambda)^{\,k-t}\,\delta_{k}},0\Bigr{)}$ (1)
where $\gamma$ is the discount factor, $\lambda$ is the GAE coefficient and $\delta_{t}=r_{t}+\gamma\,V(s_{t+1})-V(s_{t})$ is the one-step TD-error at timestep $t$. MaxMC uses the highest undiscounted return observed on the task instead of a bootstrap target. Other criteria such as policy entropy, one-step TD error, GAE, policy min-margin, and policy least-confidence have also been evaluated in curriculum learning contexts [5].
3 Method
At each step, curriculum design aims to select the tasks that will most effectively drive the agent’s learning progress. Our approach enhances UED by combining (1) a refined regret approximation integrating both value and transition errors and (2) a simple yet principled measure of task co-learnability. We compute a unified Task Priority score from these two quantities, which then governs both new task generation and replay sampling in the UED loop.
3.1 Regret Approximation via Transition Prediction Loss
Since regret quantifies the difficulty an agent experiences on a task, improving regret approximation can lead to more accurate task difficulty estimation and thus more effective curricula.
To address this, we decompose one-step regret at state–action pair $(s,a)$ into two interpretable terms:
$\mathrm{Regret}(s,a)$ $=Q^{*}(s,a)-Q(s,a)$ (2)
$=\gamma\Bigl{(}\mathbb{E}_{s_{1}\sim P(\cdot\mid s,a)}\bigl{[}Q^{*}(s_{1},a^{\prime})\bigr{]}-\mathbb{E}_{s_{2}\sim\hat{P}(\cdot\mid s,a)}\bigl{[}Q(s_{2},a^{\prime\prime})\bigr{]}\Bigr{)}$
$=\gamma\Bigl{(}\underbrace{\mathbb{E}_{s_{1}\sim P}\bigl{[}\,Q^{*}(s_{1},a^{\prime})-Q(s_{1},a^{\prime})\,\bigr{]}}_{\text{(i) Value Error}}+\underbrace{\mathbb{E}_{s_{1}\sim P}\bigl{[}Q(s_{1},a^{\prime})\bigr{]}-\mathbb{E}_{s_{2}\sim\hat{P}}\bigl{[}Q(s_{2},a^{\prime\prime})\bigr{]}}_{\text{(ii) Transition Prediction Error}}\Bigr{)}$
Term (i), the value error, measures the expected discrepancy between the true optimal action-value $Q^{*}$ and the agent’s learned $Q$. Term (ii), the transition prediction error, quantifies the bias introduced by replacing the true transition kernel $P$ with an empirical approximation $\hat{P}$, where $\hat{P}(s^{\prime}\!\mid s,a)$ is the conditional distribution estimated from the agent’s collected experience $\{(s,a,s^{\prime})\}$. Accurately estimating both terms is crucial: term (i) reflects the agent’s value-approximation quality, whereas term (ii) captures how model-mismatch in the dynamics affects long-term returns.
Because $Q^{*}$ and $P$ are unknown, we approximate (i) via the Positive Value Loss (PVL) and (ii) via a learned transition dynamics estimator $f_{\phi}$. Concretely:
Definition 1 (Average Transition Prediction Loss). Train a transition dynamics estimator $f_{\phi}$, implemented as a recurrent model, to minimize a one-step reconstruction loss $L_{\mathrm{trans}}(s_{t},a_{t})$ between the observed next state $s_{t+1}$ and the prediction $\hat{s}_{t+1}=f_{\phi}(s_{t},a_{t})$ (details in Appendix F.3). Over an episode $\tau=(s_{0},a_{0},\dots,s_{T})$, define
$\mathrm{ATPL}(\tau)=\frac{1}{T}\sum_{t=0}^{T}L_{\mathrm{trans}}(s_{t},a_{t}).$ (3)
Combine the two estimates into a single scalar:
$\widehat{\mathrm{Regret}}(\tau)=\mathrm{PVL}(\tau)\;+\;\alpha\,\mathrm{ATPL}(\tau),$ (4)
where $\alpha>0$ balances value versus transition terms (see Appendix B).
By explicitly incorporating both components, our regret approximation more faithfully reflects task difficulty and yields more effective curricula. After computing the approximated regret $\widehat{\mathrm{Regret}}(\tau)$ for each sampled task instance, we append this value to the Task Difficulty Buffer (TDB). Consequently, the number of entries in the TDB corresponding to task $i$ exactly equals the number of times task $i$ has been sampled (see Figure 2 for an overview).
Figure 2: Task Difficulty Calculation Workflow. The Task Difficulty Buffer (TDB) records each task’s history of approximated regret. The agent interacts with sampled tasks to collect episode trajectories, which are stored in the Rollouts Buffer (RB). For each trajectory, we compute the Positive Value Loss (PVL) and the Average Transition-Prediction Loss (ATPL). Their sum produces the updated task difficulty (approximated regret), which is appended to the TDB to refresh each sampled task’s stored difficulty.
3.2 Task Priority Construction
A central challenge in UED is to decide which tasks to present to the agent at each step. Intuitively, we want to (1) focus on tasks that remain challenging, while (2) exploiting tasks whose training yields transfer to others. To this end, we introduce two complementary quantities, Task Difficulty and Co-Learnability, and combine them into a single Task Priority score.
Definition 2 (Task Difficulty). Let $s_{i}(t)\;=\;\max\{\,s\leq t:\mathrm{TDB}(i,s)\text{ exists}\}$ denote the most recent time at or before $t$ when task $i$ was sampled (and its approximated regret stored). Then we define
$\mathrm{TaskDifficulty}(i,t)=\begin{cases}\mathrm{TDB}\bigl{(}i,\,s_{i}(t)\bigr{)},&\text{if }s_{i}(t)\text{ is finite},\\[4.0pt] 0,&\text{if }i\text{ has never been sampled before }t.\end{cases}$ (5)
This ensures that $\mathrm{TaskDifficulty}(i,t)$ always reflects the most recent approximated regret for task $i$, with larger values indicating greater remaining challenge.
Definition 3 (Co-Learnability). Beyond difficulty, we wish to capture how training on one task accelerates progress on others. Let $\mathcal{T}_{t+1}$ be the set of tasks replayed at time $t+1$. We define
$\mathrm{CoLearnability}(i,t)=\frac{1}{|\mathcal{T}_{t+1}|}\sum_{j\in\mathcal{T}_{t+1}}\bigl{[}\mathrm{TaskDifficulty}(j,t)-\mathrm{TaskDifficulty}(j,t+1)\bigr{]}$ (6)
i.e.,  the average reduction in difficulty of the replayed tasks induced by sampling $i$. Positive co-learnability indicates that visiting $i$ yields broad transfer benefits.
Proposition 1. Assuming $\mathcal{T}_{t+1}=\Lambda$ and defining
$Y_{j}\;=\;\mathrm{TaskDifficulty}(j,t+1)\;-\;\mathrm{TaskDifficulty}(j,t),$
the co-learnability of task $i$ at time $t$ coincides with the conditional expectation, that is
$\mathrm{CoLearnability}(i,t)=\mathbb{E}_{j\sim\mathcal{T}_{t+1}}\bigl{[}Y_{j}\mid T_{i}=1,\,t\bigr{]}.$
Proposition 1 confirms that our co-learnability definition precisely measures the mean one-step change in task difficulties, conditional on having replayed task i, when the next sampled task set covers the entire buffer.
Definition 4 (Task Priority). We combine difficulty and Co-Learnability into a scalar score, then pass it through a rank function, which replaces raw values by their relative ordering, to obtain robust sampling weights as in ACCEL [7]:
$\mathrm{TaskPriority}(i,t)=\mathrm{Rank}\Bigl{(}\mathrm{TaskDifficulty}(i,t)+\beta\,\mathrm{CoLearnability}(i,t)\Bigr{)}$ (7)
where $\beta>0$ trades off challenge versus transfer. We sample tasks for replay proportionally to their computed priority scores, so that higher-priority tasks are more likely to be selected at each step. By ranking, we prevent outliers from dominating the curriculum and ensure a smooth sampling distribution.
Theorem 2 (Expected Staleness). Assume that at each replay step $s\geq 0$, tasks in buffer $\Lambda$ are sampled according to
$\Pr\bigl{[}i\text{ at }t_{0}+s\bigr{]}=\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)}\,,$
[…]
Then the expected number of steps $T=t_{i}-t_{0}$ until task $i$ is sampled satisfies
$\mathbb{E}[T]\;\leq\;\frac{|\Lambda|\,M_{0}}{\bigl{(}\mathrm{TaskPriority}(i,t_{0})\bigr{)}}\,.$
This guarantees that no task remains “stale” indefinitely: its expected waiting time scales inversely with its initial priority. While the boundedness and non-increasing requirements in Theorem 3.2 may initially seem restrictive, they follow naturally from the agent’s continual learning process. Indeed, Figure 10 shows an overall upward trend in returns during training, indicating that the agent’s performance steadily improves and, consequently, task priority scores tend to decrease over time.
3.3 Overall UED Workflow
Figure 3: Method Workflow Overview. The three panels in figure depict: (1) Task Sampling: levels are drawn from the buffer in proportion to their priority scores. (2) Priority Update: we recompute each level’s priority based on our task priority definition (Eq. 7). (3) Task Mutation: the lowest-priority levels are mutated into new variants and reinserted into the buffer.
The overall UED algorithm follows the same loop as ACCEL [7], with the sole modification that task scoring uses our task priority definition (Eq. 7) in place of positive value loss alone. We maintain a policy $\pi_{\phi}$, a level generator G, a Task Buffer $\Lambda$, and the sampling distribution $P_{D}$ over the decision variable $d\in\{0,1\}$, where $d=0$ indicates generator sampling and $d=1$ indicates buffer replay.
At each iteration $t$, we draw $d_{t}\sim P_{D}$ and proceed as follows. If $d_{t}=0$, we sample a fresh task parameter $\theta$ from the generator and run the current policy to collect a trajectory $\tau$ on $\theta$. We then compute the approximated regret score $\widehat{\mathrm{Regret}}(\tau)$ using Eq. 4. If $\widehat{\mathrm{Regret}}(\tau)$ exceeds a predefined threshold, $\theta$ is appended to $\Lambda$.
If instead $d_{t}=1$, we sample a batch of $B$ tasks $\{\theta_{k}\}_{k=1}^{B}$ from $\Lambda$ according to their TaskPriority distribution (Eq. 7). For each $\theta_{k}$, we execute the policy $\pi_{\phi}$ to collect trajectory $\tau_{k}$, perform the standard RL update on $\phi$, and compute its approximated regret $\widehat{\mathrm{Regret}}(\tau_{k})$. We then select the $N_{\text{mutate}}$ tasks with the lowest regret scores and feed each through the task editor to obtain a mutated variant $\tilde{\theta}$. Finally, we replace the corresponding entries in $\Lambda$ with these $\tilde{\theta}$, ensuring that easier tasks are systematically evolved into more challenging ones.
Following each update to the Task Difficulty Buffer, we recompute the $\mathrm{TaskDifficulty}(i,t)$ for every modified task $i$ and update the $\mathrm{CoLearnability}(j,t-1)$ for all previously sampled tasks $j$. These refreshed scores determine the priorities for the next replay phase. The algorithm thus alternates between sampling new tasks and performing prioritized replay until the policy performance converges. The full procedure is detailed in Algorithm 1.
4 Experiments
We evaluate TRACED on two procedurally generated domains: MiniGrid (MG) and BipedalWalker (BW). In both environments, we compare against DR, PLR, ACCEL (our primary baseline, since TRACED builds directly upon it), and ADD. In BW only, we additionally include the recent sota method CENIE [20]. Detailed descriptions of all baselines can be found in Appendix M.
For each domain, we track the emergent complexity of the sampled curricula during training and evaluate zero-shot test performance on held-out levels. To summarize test results, we report the median, interquartile mean (IQM), mean, and optimality gap using the rliable library [21]. The optimality gap measures how far an algorithm falls short of a target performance level, beyond which further gains are deemed negligible. Accordingly, higher IQM and lower optimality gap values indicate better performance.
All methods use Proximal Policy Optimization (PPO) [22] as the student agent. Following prior work [7, 20], we plot the performance against the number of PPO updates; the corresponding environment interactions per update are reported in Appendix C. For more details on hyperparameters and network structures, refer to Appendix F.
SixteenRooms
Maze
Labyrinth
LargeCorridor
(a) Maze
Stairs
Roughness
(b) BipedalWalker
Figure 5: Held-out Evaluation Environments. (a) Example held-out mazes for zero-shot evaluation: 4 tasks are shown (see Appendix F.1 for all 12 task definitions). (b) Example held-out BipedalWalker terrains for zero-shot evaluation: 2 tasks are shown (see Appendix F.2 for all 6 task definitions).
4.1 Partially Observable Navigation
We evaluate our curriculum design on a partially observable maze navigation domain based on MiniGrid [16], as in prior UED work. Four representatives are shown in Figure 5. The agent observes a 147-dimensional pixel observation and is trained for up to 20k PPO updates.
Figure 6: Zero-Shot Transfer Performance in MiniGrid. Aggregated solved rates on held-out MiniGrid mazes after 10k and 20k PPO updates. TRACED at 10k updates (red) matches or outperforms DR (orange), $\text{PLR}^{\perp}$ (green), ADD (gray), and ACCEL (blue) at 20k updates. Averaged over 10 random seeds.
As shown in Figure 6, our method achieves a superior solved rate at only 10k updates, matching or exceeding the 20k update performance of all baselines. In particular, TRACED’s median solved rate at 10k surpasses ACCEL at 20k, and its IQM leads the field, indicating that the majority of runs benefit rapidly from our combined regret and co-learnability scoring. The mean performance mirrors this ordering, showing that even lower-performing seeds see substantial gains. Furthermore, TRACED at 10k produces the smallest optimality gap, closing the distance to an ideal target success rate faster than any competitor; in contrast, ACCEL requires only 20k updates to halve its gap.
Table 1: Wall-clock training time comparison. Average training duration (hours $\pm$ SE over 10 runs) on the MiniGrid domain for each algorithm at 10k and 20k PPO updates.
TRACED 10k ACCEL 10k ACCEL 20k ADD 10k ADD 20k PLR 10k PLR 20k DR 10k DR 20k
13.78 $\pm$ 0.36 12.94 $\pm$ 0.66 26.58 $\pm$ 0.76 22.48 $\pm$ 0.27 45.08 $\pm$ 0.29 14.87 $\pm$ 0.62 31.83 $\pm$ 1.36 5.82 $\pm$ 0.12 12.41 $\pm$ 0.18
Compared to ACCEL, TRACED halves the wall-clock training time while maintaining equivalent or better transfer performance (Table 1). Even relative to ACCEL at the same 10k update budget, TRACED incurs a 6% computational overhead yet delivers a 22% relative increase in mean solved rate. Taken together, these metrics demonstrate that TRACED not only accelerates learning but also delivers more consistent and reliable zero-shot transfer across diverse MiniGrid mazes. Detailed per-task zero-shot results are provided in Appendix E, and the number of environment interactions per PPO update for each method is in Appendix C.
(a) Shortest Path Length
(b) Number of Blocks
Figure 7: Emergent maze complexity metrics. Shortest path length and number of blocks both grow faster under TRACED (pink) than ACCEL (blue). This faster ramp-up indicates that our curriculum more effectively escalates difficulty in lockstep with agent learning. Averaged over 10 random seeds with shaded 95% confidence intervals.
To analyze emergent environment complexity, we track two structural metrics for each generated maze: the length of the shortest solution path and the number of obstacles (blocks), as plotted in Figure 7. Both metrics rise much more rapidly under TRACED (red) than under ACCEL (blue), showing that our priority scoring more effectively distinguishes easy versus challenging tasks to produce a steadily escalating curriculum. Notably, the ACCEL + ATPL (brown) also drives complexity upward far faster than ACCEL. It closely parallels TRACED’s gains, demonstrating that the transition-prediction component alone contributes greatly to complexity ramp-up.
(a) PerfectMazeLarge
(b) PerfectMazeXL
Figure 8: PerfectMaze Evaluation. (a), (b) Two held-out maze instances, PerfectMazeLarge and PerfectMazeXL, used for zero-shot testing. (c) Zero-shot solved rates after 10k and 20k PPO updates (mean $\pm$ std over 10 random seeds). TRACED achieves the highest 10k performance on PerfectMazeLarge and closely matches the best 20k performance on PerfectMazeXL.
To stress-test our curriculum on extremely large, procedurally generated mazes, we introduce two PerfectMaze benchmarks. PerfectMazeLarge consists of $51\times 51$ grids with a maximum episode length exceeding 5k steps, while PerfectMazeXL scales this to $100\times 100$ grids. Figure 8 shows representative levels from each variant. We evaluate zero-shot transfer performance, measuring the mean success rate over 100 episodes per seed (10 seeds total). As depicted in Figure 8, on PerfectMazeLarge, TRACED achieves the highest 10k solved rate (27% $\pm$ 23%), outperforming ACCEL’s best 20k rate (20% $\pm$ 25%) and far exceeding ADD, DR and $\text{PLR}^{\perp}$. On the even more complex PerfectMazeXL, ACCEL narrowly leads at 12% $\pm$ 28% after 20k updates, with TRACED close behind at 10% $\pm$ 14% after just 10k updates, demonstrating that TRACED scales to extremely large mazes. Detailed per-baseline results for both benchmarks are provided in Appendix E.
4.2 Walking in Challenging Terrain
We further validate our curriculum in the continuous-control BipedalWalkerHardcore environment from OpenAI Gym [23], as modified by Wang et al. (2019) [24]. This domain features a procedurally generated terrain controlled by eight parameters, terrain roughness, pit gap frequency, stump height, stair spacing, etc., that jointly determine locomotion difficulty. We consider the complete set of eight parameters in our design space. Figure 5 illustrates two representative terrains with varying stair heights and surface roughness.
Figure 9: Aggregate zero-shot performance on BipedalWalker terrains. All scores are normalized by the maximum return of 300 in the BipedalWalker domain. TRACED at 10k updates (pink) matches or exceeds ACCEL-CENIE at 20k updates (purple) across all metrics. Averaged over 5 random seeds.
(a) Basic
(b) Hardcore
(c) Stair
(d) Pitgap
(e) Stump
(f) Roughness
Figure 10: Return progression on BipedalWalker terrains. TRACED (pink) consistently outperforms DR (orange), PLR (green), ADD (gray), and ACCEL (blue). Averaged over 10 random seeds with shaded 95% confidence intervals.
Our student agent is trained for 10k PPO updates, while baselines run up to 20k updates. We evaluate zero-shot transfer on six held-out test terrains over 100 episodes each. Figure 9 reveals consistent gains across all four aggregate metrics, median, IQM, mean, and optimality gap, on the six held-out BipedalWalker terrains. Specifically, after only 10k updates, TRACED outperforms all baselines at 20k updates. Although its variance is slightly higher, TRACED’s 10k-update curve matches or exceeds ACCEL-CENIE at 20k, demonstrating both superior sample efficiency and robust generalization. Moreover, as in MG, TRACED reaches the highest performance in half the wall-clock time on BW relative to ACCEL (see Appendix C). These results confirm that our curriculum design remains effective not only in discrete grid worlds but also in complex continuous domains. Refer to Appendix E for detailed results broken down by each terrain.
As shown in Figure 10, TRACED consistently surpasses ACCEL in zero-shot returns across all terrains. In the first 5k updates, TRACED trails ADD, DR, and PLR because it builds upon ACCEL’s seed curriculum, mutating the easiest environments first, before increasing difficulty. After 5k updates, however, TRACED continues to climb steadily, while ADD, DR, and PLR plateau in a stale regime. This sustained upward trajectory demonstrates that TRACED effectively escalates task complexity in tandem with the agent’s improving capabilities, yielding superior long-term performance.
4.3 Ablation Study
Figure 11: Both ATPL and CL are important design choices. Showing per-task zero-shot solved rates for 4 variants: full TRACED (pink), ATPL only (brown), Co-Learnability only (teal), and ACCEL (blue). Averaged over 5 random seeds.
To isolate the contributions, we compare TRACED at 10k updates against three ablations (Figure 11): ATPL only (TRACED – CL), Co-Learnability only (TRACED – ATPL), and the ACCEL baseline (TRACED – ATPL – CL). TRACED (pink) encloses the largest area in the radar chart, demonstrating that both ATPL and Co-Learnability are necessary. The ATPL only variant (brown) outperforms ACCEL (blue) on 9 out of the 12 held-out mazes, confirming that incorporating transition prediction error into the regret approximation enhances the agent’s generalization ability. In contrast, the Co-Learnability only variant (teal) falls below the ACCEL baseline, indicating that Co-Learnability is effective only when paired with our regret approximation. Detailed numerical results are provided in Appendix E. Additional ablations on BipedalWalker and sensitivity analyses for $\alpha$ and $\beta$ appear in Appendix B.
5 Conclusion
In this paper, we introduced Unsupervised Environment Design (UED) method with two key components: (i) an explicit transition prediction error term for regret approximation, and (ii) a lightweight Co-Learnability metric that captures cross-task transfer effects. By integrating these into the standard generator–replay loop, TRACED produces curricula that escalate environment complexity in tandem with agent learning.
Empirically, we demonstrated on two procedurally generated domains (MiniGrid navigation and BipedalWalker) that TRACED outperforms baselines (DR, PLR, ADD, ACCEL), including sota CENIE, using only half the training updates. We showed superior zero-shot transfer success rates, faster growth in structural complexity, and scalability to extremely large mazes. Ablation studies confirmed that each component is essential: ATPL drives the primary complexity ramp-up, while co-learnability yields gains when paired with our regret estimates.
Looking forward, Co-Learnability offers a simple, computationally light mechanism for capturing inter-task influences and could be further refined via more sophisticated causal estimators or learned models. More broadly, any RL setting that relies on regret-approximation stands to benefit from incorporating transition-prediction error, providing an easy-to-implement boost in sample efficiency. We anticipate that these ideas will inspire future work on adaptive curricula, advanced editing mechanisms, and broader applications of regret-guided exploration in open-ended learning environments.
Appendix A Proof
A.1 Proof of Proposition 1
Proposition 1. Assuming $\mathcal{T}_{t+1}=\Lambda$ and defining
$Y_{j}\;=\;\mathrm{TaskDifficulty}(j,t+1)\;-\;\mathrm{TaskDifficulty}(j,t),$
the co-learnability of task $i$ at time $t$ coincides with the conditional expectation, that is
$\mathrm{CoLearnability}(i,t)=\mathbb{E}_{j\sim\mathcal{T}_{t+1}}\bigl{[}Y_{j}\mid T_{i}=1,\,t\bigr{]}.$
Proof. Since $\mathcal{T}_{t+1}=\Lambda$,
$\mathbb{E}_{j\sim\mathcal{T}_{t+1}}\bigl{[}Y_{j}\mid T_{i}=1,\,t\bigr{]}\;=\;\frac{1}{|\Lambda|}\sum_{j\in\Lambda}Y_{j}=\mathrm{CoLearnability}(i,t)$
A.2 Proof of Theorem 2
Theorem 2 (Expected Staleness). Assume that at each replay step $s\geq 0$, tasks in buffer $\Lambda$ are sampled according to
$\Pr\bigl{[}i\text{ at }t_{0}+s\bigr{]}=\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)}\,,$
[…]
Then the expected number of steps $T=t_{i}-t_{0}$ until task $i$ is sampled satisfies
$\mathbb{E}[T]\;\leq\;\frac{|\Lambda|\,M_{0}}{\bigl{(}\mathrm{TaskPriority}(i,t_{0})\bigr{)}}\,.$
Proof. Let
$q_{s}\;=\;\Pr\bigl{[}\text{choose }i\text{ at time }t_{0}+s\bigr{]}\;=\;\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)}\,,$
so that $1-q_{s}$ is the probability $i$ is not chosen at step $s$.
Since for every $j\in\Lambda$ we have $\mathrm{TaskPriority}(j,t_{0}+s)\leq M_{0})$, it follows that
$\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)\;\leq\;|\Lambda|\;M_{0}.$
Hence
$q_{s}\;=\;\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j}\mathrm{TaskPriority}(j,t_{0}+s)}\;\geq\;\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{|\Lambda|\;M_{0}}\;\geq\;\frac{\mathrm{TaskPriority}(i,t_{0})}{|\Lambda|\;M_{0}}=:\;c,$ (8)
For any nonnegative integer $k$,
$\Pr[T>k]$ $=\Pr\bigl{(}\text{task }i\text{ is not chosen in any of steps }0,1,\dots,k-1\bigr{)}$
$=\prod_{s=0}^{k-1}\bigl{(}1-q_{s}\bigr{)}\;\leq\;\prod_{s=0}^{k-1}\bigl{(}1-c\bigr{)}=\bigl{(}1-c\bigr{)}^{k}.$
where the inequality uses (8).
$\mathbb{E}[T]=\sum_{k=0}^{\infty}\Pr[T>k]\;\leq\;\sum_{k=0}^{\infty}(1-c)^{k}\;=\;\frac{1}{c}\;=\;\frac{|\Lambda|\;M_{0}}{\bigl{(}\mathrm{TaskPriority}(i,t_{0})\bigr{)}}.$
This completes the proof.
Appendix B Additional Experiments
B.1 Ablation Study of Scaling Factor
In this ablation, we keep every setting in Appendix H fixed except for one of the weight-scaling factors, $\alpha$ (ATPL weight), $\beta$ (Co-Learnability weight). Deviating from defaults on either $\alpha$ or $\beta$ reduces the solved rates in most domains as shown in Table 2. For both ignoring one of the weighted terms ($\alpha=0.0$ or $\beta=0.0$) and using excessively large weight ($\alpha=100.0$ or $\beta=100.0$), solved rates show slight improvement for some individual tasks, but statistical drops for overall performance. The default TRACED weights show the best performance compared to fixed projection baselines. This result shows that an effective balance between ATPL and Co-Learnability is needed.
Table 2: MiniGrid Ablation Solved Rates. Mean $\pm$ SE of solved rates on 12 held-out MiniGrid tasks, comparing different fixed projection Weight baselines ($\alpha=0.0$, $\alpha=100.0$, $\beta=0.0$, $\beta=100.0$) and TRACED ($\alpha=1.0$, $\beta=1.0$) over 5 seeds. For TRACED, 5 runs were randomly sampled from 10 seeds to ensure consistent evaluation. Each method is evaluated after 10k PPO updates. Bold indicates the best result per task; underline indicates the second-best.
Environment $\alpha$=0.0 $\alpha$=100.0 $\beta$=0.0 $\beta$=100.0 TRACED
16Rooms 0.81 $\pm$ 0.1 0.06 $\pm$ 0.04 0.73 $\pm$ 0.09 0.74 $\pm$ 0.17 0.79 $\pm$ 0.19
16Rooms2 0.94 $\pm$ 0.05 0.0 $\pm$ 0.0 0.28 $\pm$ 0.2 0.59 $\pm$ 0.15 0.72 $\pm$ 0.17
SimpleCrossing 0.84 $\pm$ 0.04 0.43 $\pm$ 0.08 0.86 $\pm$ 0.04 0.82 $\pm$ 0.04 0.89 $\pm$ 0.01
FourRooms 0.45 $\pm$ 0.05 0.23 $\pm$ 0.04 0.41 $\pm$ 0.05 0.52 $\pm$ 0.03 0.47 $\pm$ 0.02
SmallCorridor 0.42 $\pm$ 0.24 0.12 $\pm$ 0.06 0.62 $\pm$ 0.21 0.51 $\pm$ 0.12 0.49 $\pm$ 0.17
LargeCorridor 0.44 $\pm$ 0.26 0.04 $\pm$ 0.02 0.54 $\pm$ 0.16 0.56 $\pm$ 0.18 0.5 $\pm$ 0.14
Labyrinth 0.5 $\pm$ 0.29 0.0 $\pm$ 0.0 0.49 $\pm$ 0.25 0.19 $\pm$ 0.18 1.0 $\pm$ 0.0
Labyrinth2 0.61 $\pm$ 0.2 0.0 $\pm$ 0.0 0.2 $\pm$ 0.17 0.2 $\pm$ 0.2 0.98 $\pm$ 0.01
Maze 0.74 $\pm$ 0.21 0.0 $\pm$ 0.0 0.18 $\pm$ 0.07 0.39 $\pm$ 0.24 0.59 $\pm$ 0.17
Maze2 0.56 $\pm$ 0.26 0.0 $\pm$ 0.0 0.51 $\pm$ 0.21 0.14 $\pm$ 0.12 0.42 $\pm$ 0.19
Maze3 0.51 $\pm$ 0.27 0.0 $\pm$ 0.0 0.59 $\pm$ 0.22 0.51 $\pm$ 0.21 0.86 $\pm$ 0.07
PerfectMaze(M) 0.4 $\pm$ 0.16 0.01 $\pm$ 0.02 0.4 $\pm$ 0.06 0.34 $\pm$ 0.07 0.66 $\pm$ 0.11
Mean 0.6 $\pm$ 0.1 0.07 $\pm$ 0.01 0.5 $\pm$ 0.09 0.46 $\pm$ 0.05 0.7 $\pm$ 0.04
B.2 Ablation Study of ATPL and Co-Learnability
Table 3 breaks down zero-shot solved rates on 6 held-out BW tasks for TRACED and its ablations at 10k updates. TRACED achieves the top score in five of six tasks (Basic, Hardcore, Stairs, Stump, Roughness) and ranks second in the remaining one. The ATPL only variant (TRACED – CL) leads to the first place in one task (PitGap) and second place in two tasks (Basic, Stump), while the Co-Learnability only variant (TRACED – ATPL) only reaches the second place in two tasks (Hardcore, Roughness). Removing both components (the ACCEL baseline) ranks second on one (Stairs). These results confirm that both transition prediction error and Co-Learnability are essential to TRACED’s high and stable performance across diverse tasks.
Table 3: BipedalWalker Ablation Results. Mean $\pm$ SE of zero-shot test returns on 6 held-out BipedalWalker tasks, comparing ACCEL (10k), TRACED - ATPL (10k), TRACED - CL (10k), and TRACED (10k, $\alpha=1.5$, $\beta=0.6$) over 5 seeds. Bold indicates the best result per task; underline indicates the second-best.
Environment ACCEL 10k TRACED - ATPL 10k TRACED - CL 10k TRACED 10k
Basic 281.65 $\pm$ 5.25 282.61 $\pm$ 4.64 286.75 $\pm$ 5.01 293.67 $\pm$ 3.56
Hardcore 37.59 $\pm$ 15.0 53.77 $\pm$ 20.8 48.35 $\pm$ 23.74 86.83 $\pm$ 17.96
Stairs -38.71 $\pm$ 10.54 -43.99 $\pm$ 12.99 -34.1 $\pm$ 8.44 -29.0 $\pm$ 10.4
PitGap -65.07 $\pm$ 7.57 -72.92 $\pm$ 13.95 -5.2 $\pm$ 66.89 -39.26 $\pm$ 11.42
Stump -79.18 $\pm$ 5.45 -38.06 $\pm$ 39.88 -19.89 $\pm$ 64.11 34.16 $\pm$ 54.58
Roughness 161.72 $\pm$ 28.36 191.75 $\pm$ 25.07 191.68 $\pm$ 4.99 193.29 $\pm$ 21.6
Mean 49.67 $\pm$ 11.24 62.19 $\pm$ 11.21 77.93 $\pm$ 17.58 89.95 $\pm$ 12.95
B.3 Hyperparameter Sensitivity Analysis
Table 4 reports the impact of varying the ATPL weight $\alpha$ and the Co-Learnability weight $\beta$ in TRACED on 12 held-out MiniGrid mazes, each evaluated after 10k PPO updates. Four of the six $(\alpha,\beta)$ settings outperform the 10k update baselines (Appendix E), and even the weakest setting matches the performance of PLR at 10k. This demonstrates that TRACED delivers consistently strong zero-shot generalization across a wide range of hyperparameters, without the need for extensive tuning.
Table 4: Effect of Hyperparameter Weights $(\alpha,\beta)$ on MiniGrid Generalization. Mean $\pm$ SE of zero-shot solved rates across 12 held-out MiniGrid environments, evaluated under various $(\alpha,\beta)$ settings. Except for the final column, TRACED with $(\alpha=1.0,\beta=1.0)$, averaged over five seeds, all other entries report single-seed measurements. Bold indicates the best result per task; underline indicates the second-best.
Environment (0.75, 1.0) (1.25, 1.0) (1.0, 0.75) (1.0, 1.25) (1.25, 0.75) TRACED (1.0, 1.0)
16Rooms 0.26 0.90 1.00 0.96 1.00 $0.79\pm 0.19$
16Rooms2 0.00 0.05 0.08 0.78 0.10 $\underline{0.72}\pm 0.17$
SimpleCrossing 0.82 0.84 0.85 0.71 0.78 $\textbf{0.89}\pm 0.01$
FourRooms 0.68 0.36 0.52 0.39 0.51 $0.47\pm 0.02$
SmallCorridor 0.55 0.00 0.01 0.00 0.01 $\underline{0.49}\pm 0.17$
LargeCorridor 0.94 0.02 0.01 0.00 0.01 $\underline{0.50}\pm 0.14$
Labyrinth 0.98 1.00 1.00 1.00 1.00 $\textbf{1.00}\pm 0.00$
Labyrinth2 1.00 0.73 0.76 1.00 0.95 $\underline{0.98}\pm 0.01$
Maze 0.76 0.16 0.39 0.00 1.00 $0.59\pm 0.17$
Maze2 0.07 0.00 0.94 1.00 0.98 $0.42\pm 0.19$
Maze3 0.92 1.00 0.10 0.89 0.93 $0.86\pm 0.07$
PerfectMaze(M) 0.56 0.38 0.55 0.55 0.85 $\underline{0.66}\pm 0.11$
Mean 0.63 0.45 0.52 0.61 0.67 $\textbf{0.70}\pm 0.04$
Table 5 reports zero-shot test returns on six held-out BipedalWalker terrains after 10k PPO updates, under five different $(\alpha,\beta)$ settings. Notably, three of those configurations outperform all 10k-update baselines (Appendix E). Even the mean return across all five $(\alpha,\beta)$ choices ($\approx 64.9$) exceeds every baseline. These results demonstrate that TRACED delivers consistently strong zero-shot generalization across a wide range of weight settings, without requiring extensive hyperparameter tuning.
Table 5: Effect of Hyperparameter Weights $(\alpha,\beta)$ on BipedalWalker Generalization. Mean $\pm$ SE of zero-shot test returns on six held-out BipedalWalker terrains, for various $(\alpha,\beta)$ settings. Only the final column (TRACED with $(\alpha=1.5,\beta=0.6)$) is averaged over five seeds; all other entries are single-seed runs. Bold indicates the best result per terrain; underline indicates the second-best.
Environment (1.25, 0.6) (1.75, 0.6) (1.5, 0.45) (1.5, 0.75) TRACED (1.5, 0.6)
Basic 282.06 291.04 281.32 279.37 293.67$\pm$3.56
Hardcore –7.60 115.87 29.55 50.10 86.83$\pm$17.96
Stairs –52.78 –21.94 –42.37 –66.02 –29.0$\pm$10.4
PitGap 12.96 –52.72 –52.21 –122.54 –39.26$\pm$11.42
Stump –100.95 –40.11 –74.83 –73.93 34.16$\pm$54.58
Roughness 148.22 187.43 244.57 185.83 193.29$\pm$21.6
Mean 46.99 79.93 64.34 42.13 89.95$\pm$12.95
Appendix C Efficiency Analysis
C.1 Wall-Clock Training Time in BipedalWalker
Table 6: Wall-Clock Training Time on BipedalWalker. Average wall-clock duration (hours $\pm$ SE over five runs) for each algorithm at 10k and 20k PPO updates.
Domain TRACED 10k ACCEL 10k ACCEL 20k ADD 10k ADD 20k PLR 10k PLR 20k DR 10k DR 20k
BipedalWalker 35.64 $\pm$ 0.53 33.26 $\pm$ 1.17 70.22 $\pm$ 5.9 21.13 $\pm$ 0.11 41.58 $\pm$ 0.24 35.49 $\pm$ 1.07 71.01 $\pm$ 2.17 18.38 $\pm$ 0.1 37.03 $\pm$ 0.3
Table 6 reports wall-clock training times on the BipedalWalker domain. At 10k updates, TRACED requires $35.64\pm 0.53$ h, about 6% more than ACCEL’s $33.26\pm 1.17$h, yet delivers a 22% increase in mean solved rate (see Section 4).
C.2 Sample Complexity: Environment Steps
Table 7: Environment Steps. Total environment interactions (millions) for each method per given number of student PPO updates.
Environment PPO Updates PLR ADD ACCEL TRACED
MiniGrid 10k 82M 41M 93M 93M
MiniGrid 20k 165M 82M 185M
BipedalWalker 10k 165M 80M 174M 174M
BipedalWalker 20k 329M 160M 347M
Table 7 shows total environment steps required to achieve a fixed number of PPO updates. TRACED matches ACCEL’s sample complexity exactly, confirming that its superior generalization stems from improved curriculum design rather than additional data.
Appendix D Visualizing Curriculum Dynamics
(a) 2.5k, Low ATPL, Low Co-Learnability
(b) 2.5k, Low ATPL, High Co-Learnability
(c) 2.5k, High ATPL, Low Co-Learnability
(d) 2.5k, High ATPL, High Co-Learnability
(e) 5k, Low ATPL, Low Co-Learnability
(f) 5k, Low ATPL, High Co-Learnability
(g) 5k, High ATPL, Low Co-Learnability
(h) 5k, High ATPL, High Co-Learnability
Figure 12: Representative terrains from BipedalWalker selected by ATPL and co-learnability at two training stages. The top two rows show terrains after 2.5k PPO updates (first row: low ATPL; second row: high ATPL), and the bottom two rows after 5k updates (third row: low ATPL; fourth row: high ATPL). Within each row, the left column displays low co-learnability examples and the right column high co-learnability examples.
Figure 12 illustrates the joint impact of ATPL and Co-Learnability on terrain selection at two training milestones (2.5k and 5k PPO updates). ATPL alone captures the raw challenge level: low-ATPL terrains (first and third rows) remain relatively smooth, while high-ATPL terrains (second and fourth rows) feature larger gaps and steeper bumps. As training proceeds from 2.5k to 5k updates, the overall terrains become systematically harder, demonstrating TRACED’s ability to ramp up difficulty in lockstep with the agent’s improving skills. Co-Learnability then refines this progression by filtering out extremes: low Co-Learnability (left column) tends to generate either trivial or overwhelmingly difficult levels that can stall learning, whereas high Co-Learnability (right column) favors intermediate-difficulty terrains that transfer more effectively across tasks. Together, these visualizations confirm that ATPL drives a steadily increasing complexity trajectory, while Co-Learnability smooths it to sustain robust, transferable learning throughout the curriculum.
Appendix E Numerical Results
E.1 Minigrid Environment
Table 8: MiniGrid Zero-Shot Solved Rates. Mean $\pm$ SE of solved rates on 12 held-out MiniGrid tasks for DR, PLR, ADD, ACCEL, and TRACED, evaluated after 10k and 20k PPO updates over 10 seeds. Bold indicates the best result per task; underline indicates the second-best.
Environment Update DR PLR ADD ACCEL TRACED
16Rooms 10k 0.4 $\pm$ 0.12 0.81 $\pm$ 0.06 0.72 $\pm$ 0.1 0.76 $\pm$ 0.12 0.53 $\pm$ 0.14
20k 0.51 $\pm$ 0.13 0.86 $\pm$ 0.06 0.64 $\pm$ 0.12 0.65 $\pm$ 0.1
16Rooms2 10k 0.1 $\pm$ 0.09 0.42 $\pm$ 0.11 0.14 $\pm$ 0.05 0.48 $\pm$ 0.13 0.58 $\pm$ 0.13
20k 0.27 $\pm$ 0.11 0.55 $\pm$ 0.15 0.18 $\pm$ 0.1 0.58 $\pm$ 0.14
SimpleCrossing 10k 0.58 $\pm$ 0.05 0.72 $\pm$ 0.03 0.63 $\pm$ 0.03 0.81 $\pm$ 0.03 0.85 $\pm$ 0.03
20k 0.73 $\pm$ 0.02 0.8 $\pm$ 0.03 0.74 $\pm$ 0.04 0.85 $\pm$ 0.02
FourRooms 10k 0.32 $\pm$ 0.04 0.47 $\pm$ 0.01 0.39 $\pm$ 0.03 0.46 $\pm$ 0.02 0.46 $\pm$ 0.02
20k 0.5 $\pm$ 0.02 0.53 $\pm$ 0.03 0.61 $\pm$ 0.03 0.49 $\pm$ 0.03
SmallCorridor 10k 0.3 $\pm$ 0.07 0.52 $\pm$ 0.13 0.47 $\pm$ 0.1 0.37 $\pm$ 0.14 0.49 $\pm$ 0.13
20k 0.74 $\pm$ 0.1 0.61 $\pm$ 0.11 0.74 $\pm$ 0.04 0.22 $\pm$ 0.13
LargeCorridor 10k 0.36 $\pm$ 0.13 0.33 $\pm$ 0.13 0.35 $\pm$ 0.09 0.44 $\pm$ 0.15 0.43 $\pm$ 0.11
20k 0.54 $\pm$ 0.13 0.32 $\pm$ 0.11 0.72 $\pm$ 0.06 0.25 $\pm$ 0.12
Labyrinth 10k 0.1 $\pm$ 0.1 0.5 $\pm$ 0.15 0.22 $\pm$ 0.13 0.6 $\pm$ 0.15 0.93 $\pm$ 0.03
20k 0.58 $\pm$ 0.16 0.92 $\pm$ 0.06 0.9 $\pm$ 0.1 0.8 $\pm$ 0.13
Labyrinth2 10k 0.1 $\pm$ 0.1 0.28 $\pm$ 0.13 0.0 $\pm$ 0.0 0.52 $\pm$ 0.14 0.96 $\pm$ 0.03
20k 0.18 $\pm$ 0.12 0.64 $\pm$ 0.15 0.63 $\pm$ 0.14 0.78 $\pm$ 0.13
Maze 10k 0.02 $\pm$ 0.01 0.13 $\pm$ 0.1 0.04 $\pm$ 0.02 0.41 $\pm$ 0.13 0.7 $\pm$ 0.1
20k 0.1 $\pm$ 0.1 0.52 $\pm$ 0.16 0.2 $\pm$ 0.13 0.59 $\pm$ 0.14
Maze2 10k 0.19 $\pm$ 0.12 0.43 $\pm$ 0.14 0.09 $\pm$ 0.08 0.44 $\pm$ 0.15 0.6 $\pm$ 0.1
20k 0.16 $\pm$ 0.11 0.56 $\pm$ 0.14 0.28 $\pm$ 0.1 0.65 $\pm$ 0.13
Maze3 10k 0.27 $\pm$ 0.13 0.57 $\pm$ 0.13 0.2 $\pm$ 0.1 0.78 $\pm$ 0.13 0.81 $\pm$ 0.06
20k 0.46 $\pm$ 0.14 0.95 $\pm$ 0.04 0.54 $\pm$ 0.16 0.86 $\pm$ 0.09
PerfectMaze(M) 10k 0.06 $\pm$ 0.02 0.24 $\pm$ 0.05 0.17 $\pm$ 0.08 0.41 $\pm$ 0.05 0.64 $\pm$ 0.07
20k 0.2 $\pm$ 0.05 0.55 $\pm$ 0.08 0.36 $\pm$ 0.08 0.53 $\pm$ 0.08
Mean 10k 0.23 $\pm$ 0.05 0.45 $\pm$ 0.04 0.28 $\pm$ 0.04 0.54 $\pm$ 0.03 0.66 $\pm$ 0.03
20k 0.41 $\pm$ 0.03 0.65 $\pm$ 0.04 0.55 $\pm$ 0.04 0.6 $\pm$ 0.05
Table 9: MiniGrid Ablation Results. Mean $\pm$ SE of zero-shot solved rates on 12 held-out MiniGrid tasks, comparing ACCEL (10k), TRACED – ATPL (10k), TRACED – CL (10k), and TRACED (10k) over 5 seeds. For TRACED and ACCEL, 5 runs were randomly sampled from 10 seeds to ensure consistent evaluation. Bold indicates the best result per task; underline indicates the second-best.
Environment ACCEL 10k TRACED - ATPL 10k TRACED - CL 10k TRACED 10k
16Rooms 0.7 $\pm$ 0.19 0.72 $\pm$ 0.14 0.91 $\pm$ 0.06 0.79 $\pm$ 0.19
16Rooms2 0.44 $\pm$ 0.17 0.01 $\pm$ 0.01 0.79 $\pm$ 0.17 0.72 $\pm$ 0.17
SimpleCrossing 0.77 $\pm$ 0.02 0.68 $\pm$ 0.04 0.86 $\pm$ 0.04 0.89 $\pm$ 0.01
FourRooms 0.46 $\pm$ 0.04 0.36 $\pm$ 0.02 0.48 $\pm$ 0.04 0.47 $\pm$ 0.02
SmallCorridor 0.03 $\pm$ 0.03 0.44 $\pm$ 0.21 0.57 $\pm$ 0.2 0.49 $\pm$ 0.17
LargeCorridor 0.11 $\pm$ 0.1 0.42 $\pm$ 0.22 0.65 $\pm$ 0.21 0.5 $\pm$ 0.14
Labyrinth 0.55 $\pm$ 0.23 0.4 $\pm$ 0.24 0.49 $\pm$ 0.2 1.0 $\pm$ 0.0
Labyrinth2 0.88 $\pm$ 0.1 0.16 $\pm$ 0.11 0.62 $\pm$ 0.18 0.98 $\pm$ 0.01
Maze 0.42 $\pm$ 0.24 0.29 $\pm$ 0.19 0.63 $\pm$ 0.2 0.59 $\pm$ 0.17
Maze2 0.42 $\pm$ 0.24 0.3 $\pm$ 0.19 0.41 $\pm$ 0.15 0.42 $\pm$ 0.19
Maze3 0.8 $\pm$ 0.2 0.68 $\pm$ 0.2 0.86 $\pm$ 0.12 0.86 $\pm$ 0.07
PerfectMaze(M) 0.33 $\pm$ 0.04 0.34 $\pm$ 0.02 0.53 $\pm$ 0.07 0.66 $\pm$ 0.11
Mean 0.49 $\pm$ 0.04 0.4 $\pm$ 0.04 0.65 $\pm$ 0.03 0.7 $\pm$ 0.04
Table 10: PerfectMaze Zero-Shot Solved Rates. Mean $\pm$ SE of zero-shot success rates on PerfectMazeLarge and PerfectMazeXL after 10k and 20k PPO updates, comparing DR, PLR, ADD, ACCEL, and TRACED over 10 seeds. Bold indicates the best result per environment; underline indicates the second-best.
Environment Update DR PLR ADD ACCEL TRACED
PerfectMaze(Large) 10k 0.0 $\pm$ 0.0 0.02 $\pm$ 0.01 0.04 $\pm$ 0.04 0.09 $\pm$ 0.03 0.27 $\pm$ 0.07
20k 0.02 $\pm$ 0.01 0.19 $\pm$ 0.09 0.18 $\pm$ 0.1 0.20 $\pm$ 0.08
PerfectMaze(XL) 10k 0.0 $\pm$ 0.0 0.01 $\pm$ 0.0 0.02 $\pm$ 0.02 0.01 $\pm$ 0.01 0.1 $\pm$ 0.04
20k 0.0 $\pm$ 0.0 0.12 $\pm$ 0.09 0.14 $\pm$ 0.09 0.09 $\pm$ 0.06
Table 8 reports per-task zero-shot solved rates on the 12 held-out MiniGrid mazes, following [21]. We compare TRACED at 10k updates against DR, PLR, ACCEL, and ADD (at both 10k and 20k). TRACED at 10k achieves the highest solved rate in six mazes (16Rooms2, SimpleCrossing, Labyrinth, Labyrinth2, Maze, PerfectMaze(M)) and the second-best rate in one maze (Maze2). Averaged across all 12 mazes, TRACED attains a mean solved rate of 0.66, outperforming every other baseline configuration.
Table 9 presents zero-shot solved rates on 12 held-out MiniGrid mazes for TRACED and its ablations at 10k updates. Full TRACED attains the highest score in six mazes (SimpleCrossing, Labyrinth, Labyrinth2, Maze3, PerfectMaze(M), and FourRooms) and ranks second in the other six, with a mean solved rate of 0.70, outperforming every variant. These results confirm that both transition prediction error and Co-Learnability are critical to TRACED’s superior, consistent performance across diverse tasks.
Table 10 compares zero-shot solved rates of TRACED against DR, PLR, ADD, and ACCEL on two large-scale PerfectMaze environments: PerfectMazeLarge and PerfectMazeXL. After 10k updates, TRACED achieves the highest success rate on PerfectMazeLarge (27% ± 7%) and the third-highest on PerfectMazeXL (10% ± 4%), matching ADD at 20k. These results underscore TRACED’s ability to scale to extreme maze sizes.
E.2 BipedalWalker Environment
Table 11 compares TRACED’s zero-shot returns against DR, PLR, ADD, ACCEL, and ACCEL-CENIE on six BipedalWalker terrains. At 10k updates, TRACED attains the highest mean return in three environments (Basic, Hardcore, and Stump) and achieves an overall mean of 89.95 $\pm$ 31.72, outperforming all baselines.
Table 11: BipedalWalker Zero-Shot Test Returns. Mean $\pm$ SE of episode returns on six held-out BipedalWalkerHardcore terrains after 10k and 20k PPO updates, comparing DR, PLR, ADD, ACCEL, ACCEL-CENIE, and TRACED over 5 seeds. Bold indicates the best result per terrain; underline indicates the second-best. ACCEL-CENIE is estimated by applying the original CENIE paper’s ACCEL-to-ACCEL-CENIE performance ratio [20] to our ACCEL implementation.
Environment Update DR PLR ADD ACCEL ACCEL-CENIE TRACED
Basic 10k 112.56 $\pm$ 62.99 131.63 $\pm$ 56.58 119.43 $\pm$ 63.72 281.65 $\pm$ 5.25 273.56 $\pm$ 1.5 293.67 $\pm$ 3.56
20k 68.17 $\pm$ 44.71 97.67 $\pm$ 61.73 63.64 $\pm$ 62.71 281.85 $\pm$ 3.72 275.04 $\pm$ 3.19
Hardcore 10k -19.67 $\pm$ 7.35 -17.36 $\pm$ 5.99 3.83 $\pm$ 9.07 37.59 $\pm$ 15.0 66.83 $\pm$ 17.48 86.83 $\pm$ 17.96
20k -16.98 $\pm$ 5.8 -23.7 $\pm$ 2.05 10.48 $\pm$ 20.12 59.23 $\pm$ 25.5 84.09 $\pm$ 34.32
Stairs 10k -17.6 $\pm$ 7.21 -7.23 $\pm$ 7.16 -5.87 $\pm$ 9.07 -38.71 $\pm$ 10.54 -30.4 $\pm$ 8.35 -29.0 $\pm$ 10.4
20k -9.11 $\pm$ 7.67 -10.2 $\pm$ 4.5 -0.66 $\pm$ 9.2 -46.34 $\pm$ 5.78 -36.39 $\pm$ 11.48
PitGap 10k -34.39 $\pm$ 15.66 -46.21 $\pm$ 15.69 -17.8 $\pm$ 13.82 -65.07 $\pm$ 7.57 -39.81 $\pm$ 18.75 -39.26 $\pm$ 11.42
20k -26.03 $\pm$ 10.98 -48.72 $\pm$ 12.23 -7.65 $\pm$ 7.52 -64.89 $\pm$ 18.93 -45.92 $\pm$ 33.9
Stump 10k -24.99 $\pm$ 6.76 -27.74 $\pm$ 6.33 -26.15 $\pm$ 8.49 -79.18 $\pm$ 5.45 -60.05 $\pm$ 10.96 34.16 $\pm$ 54.58
20k -25.52 $\pm$ 12.39 -23.61 $\pm$ 2.86 3.26 $\pm$ 19.1 -67.18 $\pm$ 15.56 -46.34 $\pm$ 176.21
Roughness 10k -0.66 $\pm$ 12.01 3.55 $\pm$ 9.91 20.61 $\pm$ 13.78 161.72 $\pm$ 28.36 174.92 $\pm$ 2.25 193.29 $\pm$ 21.6
20k 0.19 $\pm$ 8.18 -2.73 $\pm$ 9.51 18.71 $\pm$ 21.59 213.48 $\pm$ 7.69 224.4 $\pm$ 8.63
Mean 10k 2.54 $\pm$ 15.44 6.1 $\pm$ 10.69 14.4 $\pm$ 14.05 49.67 $\pm$ 11.24 64.18 89.95 $\pm$ 12.95
20k -1.55 $\pm$ 8.83 -1.89 $\pm$ 8.93 14.63 $\pm$ 21.99 62.69 $\pm$8.91 75.81
Table 12 compares zero-shot solved rates across six BipedalWalker terrains for DR, PLR, ADD, ACCEL, and TRACED. A trajectory is considered “solved” if its return exceeds 230. At 10k updates, TRACED achieves the top solved rate in five of the six tasks (Basic, Hardcore, Stairs, PitGap, Stump) and ranks second on Roughness, yielding an overall mean of 0.36—substantially higher than ACCEL’s 0.29 and the other baselines. No baseline matches TRACED’s combined efficiency and reliability in zero-shot generalization.
Table 12: BipedalWalker Zero-Shot Solved Rates. Mean $\pm$ SE of solved rates, defined as the fraction of episodes exceeding a return of 230, on 6 held-out BipedalWalkerHardcore terrains (Basic, Hardcore, Stairs, PitGap, Stump, Roughness) after 10k and 20k PPO updates. Results compare DR, PLR, ADD, ACCEL, and TRACED over 5 seeds. Bold indicates the best per terrain; underline indicates the second-best.
Environment Update DR PLR ADD ACCEL TRACED
Basic 10k 0.35 $\pm$ 0.2 0.43 $\pm$ 0.21 0.34 $\pm$ 0.21 0.98 $\pm$ 0.02 1.00 $\pm$ 0.0
20k 0.16 $\pm$ 0.13 0.35 $\pm$ 0.22 0.2 $\pm$ 0.2 0.99 $\pm$ 0.0
Hardcore 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.01 $\pm$ 0.01 0.15 $\pm$ 0.03 0.28 $\pm$ 0.05
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.04 $\pm$ 0.04 0.23 $\pm$ 0.08
Stairs 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.02 0.03 $\pm$ 0.02
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.01
PitGap 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.01
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.02
Stump 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.18 $\pm$ 0.12
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.0.2 0.0 $\pm$ 0.0
Roughness 10k 0.01 $\pm$ 0.01 0.0 $\pm$ 0.01 0.01 $\pm$ 0.01 0.56 $\pm$ 0.1 0.65 $\pm$ 0.08
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.03 0.76 $\pm$ 0.03
Mean 10k 0.06 $\pm$ 0.04 0.07 $\pm$ 0.04 0.06 $\pm$ 0.04 0.29 $\pm$ 0.02 0.36 $\pm$ 0.03
20k 0.03 $\pm$ 0.02 0.06 $\pm$ 0.04 0.05 $\pm$ 0.05 0.34 $\pm$ 0.03
Appendix F Implementation Details
F.1 Minigrid Environment
SixteenRooms
SixteenRooms2
SmallCorridor
LargeCorridor
SimpleCrossing
FourRooms
Labyrinth
Labyrinth2
Maze
Maze2
Maze3
PerfectMaze
Figure 13: MiniGrid zero-shot environments used for evaluation. In SmallCorridor and LargeCorridor, the goal can appear in any of the corridor paths. SimpleCrossing and FourRooms are adapted from MiniGrid [16], and PerfectMaze from REPAIRED [6].
The partially observable navigation environment is designed as a $15\times 15$ grid maze, based on MiniGrid [16]. Each tile in the maze can be either an empty tile, a wall, a goal, or the agent itself. The empty tile is a navigable space through which the agent can move, whereas the wall is an impassable obstacle that blocks the agent’s movement. At the beginning of each episode, the initial position of the agent, the goal location, and the wall layout are randomly initialized. Up to 60 walls can be placed throughout the maze, increasing navigational complexity. At each timestep, the agent receives a $5\times 5$ local observation, along with its current direction. The action space is defined as a discrete set of seven possible actions, although only three—turn left, turn right, and move forward—are relevant to maze navigation. These actions are selected without masking the unused action outputs. A reward is provided when the agent successfully reaches the goal, calculated as $1-T/T_{\text{max}}$, where $T$ is the number of steps taken and $T_{\text{max}}=250$ denotes the maximum allowed steps per episode. If the agent fails to reach the goal within the time limit, it does not receive a reward.
To evaluate performance in partially observable navigation, we include additional results on the MiniGrid environments, covering a suite of challenging zero-shot tasks from prior UED work [4, 6]. Figure 13 displays all evaluation tasks.
F.2 BipedalWalker Environment
The environment is based on the BipedalWalkerHardcore environment of the OpenAI gym [23]. At each timestep, the agent receives a 24-dimensional observation, which includes information such as the hull angle and angular velocity, joint angles and speeds for the hips and knees, ground contact indicators for each foot, the robot’s horizontal and vertical velocities, and lidar-based distance measurements to the ground ahead. The action space is represented by a 4-dimensional vector that controls the torques applied to the robot’s two hip joints and two knee joints. Rewards are structured to encourage efficient and stable locomotion across the terrain. The agent receives positive rewards for processing while maintaining a straight hull posture, with small penalties applied for energy expenditure to discourage unnecessary motor use. The maximum achievable score is $300$. If the agent falls or moves backward, the episode ends immediately, and a penalty of $-100$ is applied. The maximum episode length is 2000 timesteps. To introduce additional difficulty and test the robustness of the agent’s locomotion policy, the environment includes a variety of challenging terrain features. These include stairs, which consist of sequences of elevated steps that the agent must ascend or descend; pit gaps, which are horizontal gaps over which the agent must jump or step over; stumps, which are vertical obstacles with varying heights and widths that obstruct movement; and surface roughness, which introduces uneven ground textures that can disrupt balance and foot placement.
We evaluate agents in the BipedalWalker-v3 environment, which features relatively smooth terrain, as well as in the more demanding BipedalWalkerHardcore-v3, which introduces a variety of challenging obstacles. The BipedalWalker-v3 focuses on basic locomotion and balance without significant terrain disturbances. Terrain in this environment does not have obstacles such as stairs, pit gaps, and stumps, and only gentle slopes are generated by adding noise in the range of approximately -1 to 1. In contrast, BipedalWalkerHardcore-v3 includes complex terrain elements such as stairs, pit gaps, stumps, and roughness, which require precise control and adaptive strategies. These obstacles are procedurally generated with randomized parameters. The stair height is randomly set to either $+1$ (ascending) or $-1$ (descending), with the number of steps and step width randomly sampled between 3–5 steps and 4–5 units, respectively. The pit width is determined by sampling a single integer between 3–5 units. Both the height and width of each stump are sampled as a single random integer between 1–3 units, ensuring the stumps are always square-shaped. Terrain roughness noise range is the same as BipedalWalker-v3.
To further probe the agent’s generalization capabilities, we include four targeted evaluation environments that isolate specific terrain challenges: Stairs, PitGap, Stump, and Roughness. Stairs specifies a fixed stair height of $2$ units, with $5$ steps and a step width between $4$ and $5$. PitGap sets the pit width to exactly $5$ units. Stump defines the stump height as $2$ units while allowing the stump width to vary between $1$ and $2$ units. Finally, Roughness generates terrain by adding noise $5$. Figure 14 displays all evaluation tasks.
(a) BipedalWalker-v3
(b) BipedalWalkerHardcore-v3
(c) Stairs
(d) PitGap
(e) Stump
(f) Roughness
Figure 14: Evaluation Terrains in BipedalWalker. Six held-out environments used to assess zero-shot generalization: (a) BipedalWalker-v3 with gentle noise-induced slopes; (b) BipedalWalkerHardcore-v3 combining multiple obstacle types; (c) Stairs, a fixed 5-step sequence requiring precise stepping; (d) PitGap, a 5-unit horizontal gap; (e) Stump, 2-unit tall obstacles with variable width; and (f) Roughness, continuous uneven ground generated by noise.
F.3 Network Structures
For partially observable navigation, the LSTM-based transition prediction model $f_{\phi}$ consists of three components: 1. an image encoder that compresses the input observation $s_{t}$ into a 128-dimensional latent embedding, 2. an LSTM module (hidden size 128, 1 layer) that processes the concatenation of this embedding and the action $a_{t}$, 3. and an image decoder that reconstructs the predicted next observation $\hat{s}_{t+1}$ from the LSTM output. We adopt an LSTM here because recurrent architectures have proven effective at capturing temporal dependencies in transition and video-prediction tasks under partial observability [25, 26].
For walking in challenging terrain (BipedalWalker), the model instead concatenates the raw state vector $s_{t}$ with $a_{t}$ and feeds this into an LSTM (hidden size 128, 1 layer); the recurrent output is then linearly mapped to the predicted next state $\hat{s}_{t+1}$.
Its instantaneous error is defined by
$\ell_{t}=\begin{cases}\|\,\hat{s}_{t+1}-s_{t+1}\|_{1},&\text{in MiniGrid environments},\\ \|\hat{s}_{t+1}-s_{t+1}\|_{2}^{2},&\text{in BipedalWalker}.\end{cases}$
We use L1 loss for MiniGrid because it preserves sharp edges and reduces blurring in image reconstruction [27], while mean squared error is a standard, reliable choice for continuous state regression in locomotion domains [28].
The PPO network serves as the student agent in our system, directly learning to optimize both the policy and value functions through interaction with the environment. The architecture follows the standard actor-critic structure where the actor head outputs an action distribution, and the critic head predicts the scalar value $V(s)$.
For partially observable navigation, the input consists of a 3-channel $5\times 5$ local image observation and a directional scalar with dimension 4 indicating the agent’s current orientation. The image is processed by a convolutional encoder with a filter size of 16, a kernel size of 3, and Rectified Linear Unit (ReLU) activations, followed by flattening. The directional scalar is passed through a fully connected layer of size 5, and the resulting embedding is concatenated with the flattened image features. The combined feature vector is fed separately to the actor and critic heads, each consisting of fully connected layers of size [32, 32]. The actor outputs a categorical distribution over discrete actions, while the critic produces a scalar value estimate $V(s)$.
For walking in a challenging terrain (BipedalWalker), the input is a flat state vector of dimension 24. This input is processed through a Multilayer Perceptron (MLP) with two hidden layers of 64 dimensions, followed by actor and critic heads. The actor head outputs the mean and standard deviation parameters of a diagonal Gaussian distribution over a continuous 4-dimensional action space, and the critic head outputs a scalar state value $V(s)$. The PPO learning process optimizes a clipped surrogate policy loss with clip parameter $\epsilon$ = 0.2 combined with a value function loss.
Appendix G Overall Workflow
Algorithm 1 summarizes our overall UED procedure under the TRACED framework. At each iteration, with probability $1-p_{\mathrm{replay}}$ the teacher enters the exploration phase by sampling a fresh level $\theta\sim\mathcal{G}$, executing the current policy $\pi_{\phi}$ to collect a trajectory, and computing its approximated regret
$\widehat{\mathrm{Regret}}(\tau)\;=\;\mathrm{PVL}(\tau)\;+\;\alpha\,\mathrm{ATPL}(\tau).$
If the buffer is full, we evict the lowest-priority task before appending. Otherwise, with probability $p_{\mathrm{replay}}$ the teacher enters the replay+mutation phase, drawing a batch of $B$ tasks from the buffer proportional to their TaskPriority scores, updating the policy on each, and recomputing regret. We then select the $N_{\mathrm{mutate}}$ tasks with the smallest regret, mutate each via the editor, evaluate the new variant, and replace its parent in the buffer. Finally, for every modified task we recompute TaskDifficulty, CoLearnability, and TaskPriority (Eq. (7)) before proceeding to the next timestep.
Algorithm 1 UED Workflow with Transition-Prediction Loss and Co-Learnability
1:Given: policy $\pi_{\phi}$, level generator $\mathcal{G}$, buffer capacity $K$, replay probability $p_{\mathrm{replay}}$, batch size $B$, mutation count $N_{\text{mutate}}$, scaling factors $\alpha,\beta$
2:Initialize Task Buffer $\Lambda\leftarrow\emptyset$, timestep $t\leftarrow 0$
3:while policy has not converged do
4:     Sample decision $d_{t}\sim\mathrm{Bernoulli}(p_{\mathrm{replay}})$
5:     if $d_{t}=0$ $\triangleright$ Exploration phase
6:         Sample new task $\theta\sim\mathcal{G}$
7:         Collect trajectory $\tau$ by executing $\pi_{\phi}$ on $\theta$ (no gradient)
8:         Compute regret estimate
$\widehat{\mathrm{Regret}}(\tau)=\mathrm{PVL}(\tau)+\alpha\,\mathrm{ATPL}(\tau)$
9:         Append $\theta$ to buffer $\Lambda$ (if $|\Lambda|>K$, remove lowest-priority)
10:     $\triangleright$ Replay + Mutation phase
11:         Sample a batch $\{\theta_{k}\}_{k=1}^{B}$ from $\Lambda$ w.p. $\propto$ $\mathrm{TaskPriority}(\cdot,t)$
12:         for $k=1\ldots B$ do
13:              Collect trajectory $\tau_{k}$ by executing $\pi_{\phi}$ on $\theta_{k}$
14:              Update policy $\phi$ using rewards from $\tau_{k}$
15:              Compute $\widehat{\mathrm{Regret}}(\tau_{k})$
16:         end for
17:         Select the $N$ tasks $\{\theta_{k^{\prime}}\}$ with smallest $\widehat{\mathrm{Regret}}(\tau_{k^{\prime}})$
18:         for each selected $\theta_{k^{\prime}}$ do
19:              $\tilde{\theta}\leftarrow$ editor($\theta_{k^{\prime}}$)
20:              Collect trajectory $\tilde{\tau}$ on $\tilde{\theta}$ (no gradient)
21:              Compute $\widehat{\mathrm{Regret}}(\tilde{\tau})$
22:              Replace $\theta_{k^{\prime}}$ in $\Lambda$ with $\tilde{\theta}$
23:         end for
24:     end if
25:     Buffer update:
26:     for each task $i\in\Lambda$ updated at $t$ do
27:         Recompute $\mathrm{TaskDifficulty}(i,t)$
28:         Recompute $\mathrm{CoLearnability}(i,t-1)$
29:         Update $\mathrm{TaskPriority}(i,t)$ via Eq. (7)
30:     end for
31:     $t\leftarrow t+1$
32:end while
Appendix H Hyperparameters
Table 13 contains the hyperparameters we use. Hyperparameter values were adopted directly from the DCD repository [29]. We perform hyperparameter tuning in MiniGrid and Bipedal environments, focusing on two key components, $\alpha$ and $\beta$. We explored values in the ranges $\alpha\in[0.5,1.0,1.5]$, $\beta\in[0.6,0.8,1.0]$.
Table 13: Hyperparameters used for training each method in each environment.
HyperParameter MiniGrid BipedalWalker
PPO
$\gamma$ 0.995 0.99
$\lambda_{\text{GAE}}$ 0.95 0.9
PPO rollout length 256 2048
PPO epochs 5 5
PPO minibatches per epoch 1 32
PPO clip range 0.2 0.2
PPO number of workers 16 4
Adam learning rate 1e-4 3e-4
Adam $\epsilon$ 1e-5 1e-5
PPO max gradient norm 0.5 0.5
PPO value clipping True False
Return normalization False True
Value loss coefficient 0.5 0.5
Student entropy coefficient 0.0 1e-3
Generator entropy coefficient 0.0 1e-2
TRACED
ATPL weight, $\alpha$ 1.0 1.5
Co-Learnability weight, $\beta$ 1.0 0.6
ACCEL
Replay rate, $p_{\text{replay}}$ 0.8 0.9
Buffer size, $K$ 4000 1000
Number of edits 5 3
Batch Size, $B$ 4 4
Number of mutated tasks, $N_{\text{mutate}}$ 4 4
Temperature 0.3 0.1
Staleness coefficient, $\rho$ 0.3 0.5
Appendix I Level Evolution
I.1 Visualization of Level Evolution in Minigrid
The visualization of how the Minigrid environment evolves as the number of blocks increases is shown in Figure 15. Each step of the evolutionary process produces an edited level that has a high learning efficiency.
Figure 15: Visualization of a single level’s evolving progression in the MiniGrid environment. Starting from top-left, ending bottom-right. This progress is automatically designed by our TRACED algorithm.
I.2 Evolution of Levels in BipedalWalker
Figure 16 shows the complexity metric results trained by 3 methods, TRACED (Ours), ACCEL + ATPL, ACCEL. Starting with plain terrain (near zero point), all 3 methods guide levels with more complex terrain. With aspect of (a) Stump height, (b) Stump height high, (e) Stair height step metrics, the results show TRACED and ACCEL + ATPL quickly evolve levels compared to ACCEL. This rapid level increasing brings, short wall-clock relative to performance and high performance results in various test environments.
(a) Stump height low
(b) Stump height high
(c) Stair height low
(d) Stair height high
(e) Stair height step
(f) Pit gap low
(g) Pit gap high
(h) Roughness
Figure 16: Emergent BipedalWalker terrain complexity metrics. Aspect of (a) Stump height low, (b) Stump height high, and (e) Stair height step, complexity grows faster under TRACED (Ours, pink) or ACCEL + ATPL (brown) than under ACCEL (blue). The result is averaged over 5 random seeds. This faster ramp-up indicates that our curriculum more effectively escalates difficulty in lockstep with agent learning.
I.3 Visualization of Level Evolution in BipedalWalker
Figure 17: Visualization of the level evolving progression in the BipedalWalker environment. Starting from top-left, ending bottom-right. in this example, starting with plain terrain, the pits are created and their number increases, then the roughness increases, then stairs and stumps are created and their number, width, and height increase. This progress is automatically designed by our TRACED algorithm.
Figure 17 shows the visualization of how the environment evolves as the complexity of the task increases. As we can see, the process starts from a plain terrain and gradually evolves to a level of terrain with increasingly complex parameters.
Appendix J Agent Trajectory Visualizations Across Environments
J.1 MiniGrid
Figure 18 shows a representative rollout in the PerfectMaze environment.
Figure 18: Agent trajectory visualization in the PerfectMaze evaluation task. TRACED enables efficient path planning and robust generalization to challenging maze environments, as demonstrated by the agent’s successful navigation to the goal (green).
J.2 BipedalWalker
Figure 19 illustrates a zero-shot rollout in BipedalWalkerHardcore. The agent dynamically modulates joint torques to ascend and descend steps, leap over gaps, step onto stumps, and maintain balance on uneven ground, confirming TRACED’s effectiveness in teaching transferrable locomotion skills.
Figure 19: Agent trajectory visualization in the BipedalWalkerHardcore. TRACED agent successfully overcomes various combined obstacles, including stairs, pit gaps, stumps, and roughness, demonstrating robust generalization to complex environments.
Appendix K Difference with Former Studies
K.1 Difference with ACCEL
Our approach builds on ACCEL’s evolutionary, regret-based curriculum design, refining both its regret estimator and its sampling criterion. Like ACCEL, we maintain a level buffer and record each task’s difficulty. However, instead of relying solely on Positive Value Loss (PVL) to approximate regret, we decompose regret into value error and transition-prediction error and incorporate both components into our difficulty measure.
Furthermore, prior UED methods sample levels independently based solely on their individual difficulty, ignoring cross-task transfer effects. We address this with our Co-Learnability metric, which quantifies how training on one task influences regret on others by measuring the average change in their difficulty after replay. By integrating Co-Learnability into our priority score, we can prioritize tasks that not only challenge the agent but also deliver broad transfer benefits across the entire task space.
K.2 Difference with CENIE
CENIE introduces a complementary curriculum strategy by explicitly measuring environmental novelty, rather than relying solely on regret signals as in PLR or ACCEL. It fits a Gaussian mixture model (GMM) to the agent’s past trajectories and scores new tasks by how poorly the GMM explains them, thereby steering the agent toward unfamiliar regions of the state–action space. This model-agnostic approach can be applied on top of any UED framework, including PLR and ACCEL.
However, CENIE’s diversity-driven objective differs from ours. TRACED’s primary goal is to approximate task difficulty more faithfully, by decomposing regret into value and transition-prediction errors, and to capture how training on one task transfers to others via our Co-Learnability metric. While CENIE accounts for per-task novelty and regret, it ignores cross-task transfer effects. Moreover, TRACED is agnostic to the choice of level generator or editor, meaning our method could be combined with CENIE’s novelty scoring to yield a curriculum that balances both transfer-aware difficulty and environmental diversity.
Appendix L Related Work
Unsupervised Environment Design (UED) defines a co-evolutionary framework in which a teacher agent generates task instances for a student policy, selecting those tasks with high learning potential to promote robust generalization to unseen environments [24, 30, 4, 19]. PAIRED [4] trains the teacher to maximize the regret between a protagonist and antagonist policy, causing the teacher to create increasingly challenging yet solvable environments. However, non-stationary task distributions and the high dimensionality of the task space impede convergence to optimal curricula. To mitigate these limitations, Mediratta et al. [9] introduce entropy bonuses, alternative optimization algorithms, and online behavioral cloning, and Clutr [31] employs variational autoencoder–based unsupervised task representation learning to compress the task manifold and yield more stable curricula.
Another line of work is based on Prioritized Level Replay (PLR) [5], which avoids training a teacher agent and instead randomly generates tasks while introducing a replay buffer that stores previously generated tasks. This approach prioritizes tasks with high learning potential from the buffer to more effectively train the student agent. PLR  [6] make improvement of PLR, which used stop-gradient on trajectories from newly generated tasks but updated only with replayed tasks from buffer. This leads to better performance on unseen environments counterintuitively. ACCEL [7] extends the PLR by implementing an evolutionary approach that makes small mutation to high-regret levels from the buffer. This evolutionary mechanism enables the development of progressively more complex challenges starting with simpler tasks. However, these methods [5, 6, 7] approximate regret solely via value loss, thereby implicitly learn true environment dynamics that could improve regret estimation.
Moving beyond regret-based approaches, CENIE [20] integrates novelty with regret to expose agents to more diverse learning situations. Using Gaussian Mixture Models, CENIE quantifies how much a new task’s trajectory differs from previous experiences stored in the buffer. This curriculum-awareness leads to better exploration and generalization. DIPLR [32] introduces a diversity that quantifies the similarity between different levels by computing the Wasserstein distance between their occupancy distributions over state–action trajectories. MBeDED [33] introduces the marginal benefit metric to quantify the performance gain of a student policy from training on a generated task. It compares a base policy (before training) and the updated student policy by measuring the difference in their expected returns on that task. So far, to the best of our knowledge, there are no prior works that explicitly consider Co-Learnability between tasks.
UED’s benefit on robust generalization performance in unseen environments has led to its application in other domains. For Multi-Agent Reinforcement Learning, MAESTRO [34] extends UED to competitive multi-agent settings by creating a regret-based curriculum over the joint environment and co-player space, leveraging self-play to generate opponents while training agents robust to environment and co-player variations. In contrast, RACCOON [10] applies UED to cooperative scenarios, developing a curriculum that prioritizes high-regret partners from pre-trained partners to enhance collaboration. For meta-RL, GROOVE [35] suggested policy meta-optimization (PMO) introducing algorithmic regret (AR), which measures the performance gap between a meta-learned optimizer and RL algorithms like A2C, creating curricula that identify informative environments for meta-training. DIVA [36] is the first method that extends UED to semi-supervised environment design, employing quality-diversity search to maintain a population of diverse training tasks in open-ended simulators, maximizing behavioral coverage. In sim-to-real settings, GCL [37] aligns the simulated curriculum with actual deployment tasks by sampling from real-world task distributions and adapting subsequent task generation based on the robot’s past performance.
Similar to UED works, Curriculum Reinforcement Learning (CRL) accelerates agent training by constructing a sequence of progressively challenging tasks rather than directly confronting the agent with a single, complex target task [38]. Early approaches distinguished between fixed curricula in which task order is predefined and self-paced learning, in which task selection adapts dynamically based on the agent’s learning progress [39]. Modern CRL methods have advanced this paradigm by framing curriculum generation as an inference process [40, 41]. WAKER [42] leverages reward-free curricula to train robust world models, thereby underscoring the versatility of self-paced approaches in diverse learning scenarios. However, these approaches assume access to the target task distribution, which conflicts with UED’s core assumption of an unknown distribution.
Appendix M More Details On Baseline Algorithms
In Unsupervised Environment Design (UED), Domain Randomization (DR) constructs the curriculum by uniformly sampling tasks from the environment parameter space $\Theta$. Formally, DR samples tasks according to $\theta\sim p(\Theta)$ where $p(\Theta)$ represents a uniform distribution over the task parameter space. The agent is trained on these randomly sampled tasks.
Robust PLR (PLR[6] was introduced as an extension of the original PLR [5]. While DR treats all tasks equally, PLR prioritizes tasks with high learning potential, maintaining a buffer $\Lambda$ of previously encountered tasks for replay. When sampling new tasks, the agent’s policy parameters are stop-gradiented, meaning the policy is not updated on trajectories collected from these newly sampled tasks. The collected trajectories are used to compute PVL scores to decide whether to add the task to the replay buffer. Policy updates are exclusively performed on tasks sampled from the replay buffer. By stopping the gradient on randomly sampled levels, PLR ensures that the policy is only updated on tasks specifically selected to maximize regret, leading to more robust generalization compared to both DR and PLR unintuitively.
Adversarially Compounding Complexity by Editing Levels (ACCEL) [7] actively evolves environments through an editing mechanism, allowing it to more efficiently explore the environment design space. This enables reusing the structure of sampled levels in the buffer for high-regret, rather than curating randomly sampled levels for high-regret. ACCEL maintains a buffer $\Lambda$ and employs a cycle of sampling, editing, and curation. The key insight of ACCEL is that regret serves as a domain-agnostic fitness function for evolution, enabling it to produce batches of levels at the frontier of agent capabilities. The editing mechanism involves making small mutations to previously high-regret levels, which can operate directly on environment elements such as blocks in MiniGrid. This evolutionary process creates an expanding frontier that matches the agent’s capabilities, starting with simple levels and progressively increasing in complexity. Like PLR, ACCEL employs stop-gradient when evaluating new or edited levels to ensure theoretical guarantees.
Coverage-based Evaluation of Novelty In Environment (CENIE) [20] introduces environment novelty as a complementary objective to regret in UED. CENIE quantifies environment novelty through state-action space coverage derived from the agent’s accumulated experiences across previous environments in its curriculum. The framework operates on the intuition that a novel environment should induce unfamiliar experiences, pushing the student agent into unexplored regions of the state-action space. At the core of CENIE’s implementation is the use of Gaussian Mixture Models (GMMs) to model the distribution of state-action pairs from the agent’s past experiences. Given a state-action buffer $\Gamma$ containing pairs collected from previous environments, CENIE fits a GMM with parameters $\lambda_{\Gamma}$ to represent this distribution. The novelty of a candidate environment is then quantified by measuring the similarity between its newly observed state–action pairs and the learned distribution of past state–action experiences. CENIE integrates with existing UED algorithms by combining both novelty and regret into a unified priority score for environment selection.
Adversarial Environment Design via Regret-Guided Diffusion Models (ADD) [19] frames environment generation as a diffusion process steered by the agent’s regret. Starting from uniformly sampled tasks, a pretrained score network $s_{\phi}$ is guided by the gradient of a differentiable regret estimate, computed via a learned return critic, during reverse diffusion.
In our empirical study, DR, PLR, and ACCEL were evaluated with the DCD codebase [29], and assessed ADD with the ADD codebase [19]. CENIE’s performance data were incorporated from its original publication, as no open-source implementation was available.
Appendix N Long-term Analysis on TRACED
Maze
FourRooms
Labyrinth
PerfectMaze(M)
SimpleCrossing
SmallCorridor
Figure 20: TRACED Solved-Rate Time Series. Solved rate progression on six held-out MiniGrid tasks plotted over 0–45k PPO updates for a single random seed. Each curve shows the instantaneous solved rate at each update step.
Maze
FourRooms
Labyrinth
PerfectMaze(M)
SimpleCrossing
SmallCorridor
Figure 21: ACCEL Solved-Rate Time Series. Solved rate progression on six held-out MiniGrid tasks plotted over 0–45k PPO updates for a single random seed. Each curve shows the instantaneous solved rate at each update step.
TRACED reaches near-peak solved rates very quickly. However, its performance subsequently oscillates rather than remaining perfectly stable (Figure 20). Crucially, this same post-peak fluctuation also appears under the ACCEL curriculum (Figure 21), where the agent simply arrives more slowly at its maximum but then exhibits the identical up-and-down pattern. In other words, these oscillations reflect an inherent tendency of the RL policy to suffer occasional catastrophic failures in held-out levels, rather than a pathology introduced by our task-prioritization strategy. Addressing and damping these fluctuations in the student agent’s performance is an important direction for future work.
Appendix O Limitations
As with any research, our approach presents several limitations that highlight opportunities for future investigation:
  • Co-Learnability In this paper, we proposed Co-Learnability, simple yet effective method. Follow-up studies could address these gaps by developing multi-step Co-Learnability estimators, incorporating importance-weighting or counterfactual techniques to reduce sampling bias, and combining observational co-learn signals with auxiliary models (e.g., task-conditioned value predictors) for stronger causal inference.
  • Sampling levels We sampled a fixed number of the lowest-regret levels as easy tasks to mirror ACCEL’s approach and enable fair comparisons with baseline methods. However, this simple strategy leaves room to explore more advanced task-selection mechanisms.
  • Heuristic weighting We introduced weighted approach to regret approximation and task prioritization that relies on two key parameters $\alpha$, PVL and ATPL in the regret approximation, and $\beta$, which weights the relative importance of Co-Learnability against task difficulty. While these parameters were manually tuned in the current implementation and demonstrated effective results across environments, future work could explore more principled methods for automatically determining these weights.
  • Level mutation We adopted the same approach as ACCEL which employs random mutations. Our approach focused on improving task priority rather than advancing environment parameter mutation mechanisms. The current random mutation process creates a clear opportunity for future work to advance these methods using evolutionary algorithms such as CMA-ES [43], so the Editor can offer the agent more informative tasks.
  • Extension to other RL algorithms Our experiments have been conducted with PPO-based student agents. While this allows for fair comparisons with previous work, there is a much broader range of algorithms that may interact differently with our regret-based curriculum. Off-policy methods such as SAC [44] and TD3 [45] feature distinct exploration strategies, which may yield different performance dynamics under our curriculum framework. Moreover, Model-based approaches like DreamerV3 [46] build world models for planning, offering complementary insights into curriculum design. Evaluating these methods under our curriculum framework remains an important direction for future work.
  • Extend to other UED methods We empirically demonstrated that TRACED yields more effective curricula, thereby enhancing agents’ generalization capabilities at half the training cost. While our experiments focus on the ACCEL framework, future work will integrate our approach into alternative UED methods (PLR [6] and CENIE [20]) and conduct comprehensive experiments to assess its performance in these settings.
  • Evaluation in other domains Our experimental validation has been conducted in MiniGrid [16, 4] and BipedalWalker [23, 7]. Evaluating TRACED in other domains such as Car Racing [23, 6], physical robotics applications [47], and reasoning task ARC [48] would provide a more comprehensive assessment of the approach’s generality and scalability across diverse settings.
Appendix P Experimental Setup and Reproducibility
All experiments were conducted using Python 3.8.20 on Ubuntu 22.04.4 LTS. The hardware setup included an AMD EPYC 7543 32-core processor, 80GB of RAM, and an NVIDIA A100 GPU. Experiments were run on a shared server with 8 GPUs, with each experiment using a single GPU. We implemented all experiments in Python 3.8.20, using PyTorch(v1.9.0+cu111). The environment stack was built using MiniGrid (v1.0.1), Gym (v0.15.7).